# Dynamics of non-Hermitian Floquet Wannier-Stark system

### v1

@@ V Dynamical exponents
where $|n\rangle$ denotes the eigenstates of the instantaneous Hamiltonian. The Hamiltonian $H_{\mathrm{eq}}$ can be diagonalized in the form
@@
$\left|\Psi(t)\right\rangle$ $=$ $\sum\limits_{l}a_{l}(t)|l\rangle$ (28)
$=$ $\sum\limits_{l}a_{l}(0)\sum_{n}i^{n-l}J_{n-l}(-\kappa_{0}t)|n\rangle,$
@@
Figure 3: The same plots of $P_{n}(t)$ and $P(t)$ as those in Fig. 2 are shown for several typical values of system parameters but at a resonance $\mathcal{\omega}_{0}=\omega=1$. Here, (a), (b), and (c) correspond to the cases with $\kappa_{0}=1$, $i$, and $e^{i\pi/4}$, respectively. The middle panels correspond to the probability distributions at moments $t=3T$, $6T$, and $9T$, respectively. The right panels are the logarithmic plots of total probabilities as functions of time. These results accord with our analytical analysis in the main text.
@@
The feature of the Bessel function tells us that there exists a maximum at the edge of $\mathcal{P}_{n}(t)$, which can be regarded as the wave front of the occupied-energy-level spreading. The location of such a wave front can then be determined by the equation
$\frac{\partial\mathcal{P}_{m_{c}}(t)}{\partial t}=0.$ (32)
Based on the relations $J_{n-1}(x)-J_{n+1}(x)=2J_{n}^{\prime}(x)$ and $J_{m_{c}+1}\left(-t\right)+J_{m_{c}-1}\left(-t\right)=2\left|m_{c}/t\right|J_{m_{c}}\left(-t\right),$ we have
$2J_{m_{c}+1}(-t)=2|m_{c}/t|J_{m_{c}}(-t).$ (33)
For large $n$ and $t$, we have $J_{m_{c}+1}(-t)\approx J_{m_{c}}(-t)$, which results in
$|m_{c}|\approx t.$ (34)
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1$. (ii) $\kappa_{0}=i$; in parallel, the corresponding amplitude is the Bessel function with imaginary argument
$a_{2n}(t)$ $=$ $(-1)^{n}J_{2n}(-it)$ (35)
$=$ $\frac{(-1)^{n}}{2\pi}\int_{-\pi}^{\pi}\cos(2nk)e^{-t\sin k}\mathrm{d}k,$
@@
$a_{2n}(t)\approx e^{-\frac{(2n)^{2}}{t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-t\sin k}\mathrm{d}k,$ (37)
@@
$a_{2n+1}(t)\approx e^{-\frac{(2n+1)^{2}}{t}}\frac{1}{2\pi i}\int_{-\pi}^{\pi}e^{-t\sin k}\mathrm{d}k,$ (38)
@@
$\mathcal{P}_{n}(t)=|a_{n}(t)|^{2}\approx\sqrt{\frac{2}{\pi t}}\mathcal{P}e^{-\frac{2n^{2}}{t}},$ (39)
@@
Obviously, the profile of $\mathcal{P}_{n}$ is a Gaussian type, and the total probability grows exponentially as $e^{2t}$ approaches the large $t$ limit. In this situation, the dynamics of occupied-energy-level spreading can be characterized by the full width at half maximum (FWHM),
@@
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1/2$. To verify and demonstrate the above analysis, numerical simulations are performed to investigate the dynamic behavior of dynamic processes at resonance with $\mathcal{\omega}=\mathcal{\omega}_{0}$. For the same initial state as shown in Fig. 2, the plots of $P_{n}(t)$ and $P(t)$ for several typical values of $\kappa_{0}$  are presented in Fig. 3. These numerical results accord with our above analysis in two aspects: (i) At resonance, the dynamics are no longer periodic; (ii) The level spreading exhibits distinct behaviors for Hermitian and non-Hermitian systems.
VI 2D simulator via wavepacket dynamics
In this section, we will show how the time evolution of a wavepacket in an engineered time-independent 2D square lattice can simulate our results for Floquet dynamics in a 1D time-dependent lattice. In experiments, single-particle hopping dynamics can be simulated by discretized spatial light transport in an engineered 2D square lattice of evanescently coupled optical waveguides Christodoulides2003 . A 2D lattice can be fabricated by coupled waveguides, by which the temporal evolution of the single-particle probability distribution in the 2D lattice can be visualized by the spatial propagation of the light intensity. We consider a single-particle Hamiltonian on a square lattice
$H_{\mathrm{2D}}$ $=$ $\kappa_{0}\sum_{n,m=-\infty}^{\infty}\cos\left(qm\right)|n,m\rangle\langle n+1,m|$ (43)
$-J\sum_{n,m=-\infty}^{\infty}|n,m\rangle\langle n,m+1|+\mathrm{h.c.}$
@@
which is an array of coupled 1D Bloch systems with $q$-dependent hopping strength $\kappa_{0}\cos\left(qm\right)$. Fig. 1(b) shows a schematic of the system. According to the above analysis, the system is governed by Bloch dynamics with frequency $\omega_{0}$ for vanishing $J$, regardless of whether $\kappa_{0}$ is real or complex.
Figure 4: For numerical simulations of the dynamics, the trajectories of wavepackets evolving in the in a 2D lattice with four typical configurations were obtained. The traces of the wavepacket $P(n,m)$ for the time evolutions of the four cases are shown in (a1), (b1), (c1), and (d1). Probability distributions $p(n,m,t)$ for the time evolutions of four representative cases at $t=0$, $t=2\pi/J$, $t=4\pi/J$, $t=6\pi/J$ and $t=8\pi/J\,$are shown in (a2), (b2), (c2), and (d2), respectively. The initial states and system parameters are given in the main text. The size of the system is $60\times 30$. The energy scale of the Hamiltonian is taken as $J=1$. The probability distributions on the square lattice exhibit distinct patterns corresponding to the dynamic behaviors of Bloch breathing, oscillation, diffusion and superdiffusion.
@@
$i\frac{\partial}{\partial t}|\psi\rangle=H|\psi\rangle,$ (44)
@@
where $f\left(k_{x},t\right)$ obeys the following equation:
@@
In the case with $k_{y}\approx\pi/2$, we have $\cos k_{y}\approx k_{y}-\pi/2$. One can construct a solution by the superposition of states $|k_{x},k_{y}\rangle$ with $k_{y}\approx\pi/2$ in the form Kim2006
$\left|g\left(k_{x},t\right)\right\rangle$ $=$ $(\frac{2\alpha^{2}}{\pi})^{1/4}\int_{-\pi}^{\pi}e^{-\alpha^{2}\left(k_{y}-\pi/2\right)^{2}}e^{it2J\cos k_{y}}$ (49)
×$\times f\left(k_{x},t\right)\mathrm{d}k_{y}|k_{x},k_{y}\rangle$
$\approx$ $(\frac{1}{2\pi\alpha^{2}})^{1/4}\sum_{m}e^{im\pi/2}e^{-(1/4)(m-2Jt)^{2}/\alpha^{2}}$
×$\times f\left(k_{x},t\right)|k_{x},m\rangle,$
which is a traveling Gaussian wavepacket in the $y$ direction. It is presumed that such a solution still holds true by taking $\kappa_{0}\rightarrow\kappa_{0}\cos\left(qm\right)$ when $\kappa_{0}\cos\left(qm\right)$ varies slowly during the extension of the wavepacket in the y direction due to the locality of the solution. In this sense, the traveling wavepacket experiences an adiabatic process, after which the independence of the dynamics in both directions is maintained. Obviously, a small value of q should be needed. On the other hand, the value of $\omega_{0}$ must match $q$ when the system is employed to simulate the dynamics at the resonance. To this end, a small value of $\kappa_{0}$ should be taken to maintain the Stark ladder. Actually, under the conditions $\kappa_{0}\ll J$ and $k_{y}\approx\pi/2$, such a 2D system is equivalent to a coupled waveguide array due to the linear dispersion in the $y$ direction, in which the tunneling rate between neighboring waveguides is $m$ dependent. The speed of the propagating wavepacket along $m$ is $2J$. To simulate the dynamics along $x$, one can simply take $q=\omega/\left(2J\right)$, regarding $m$ as the temporal coordinate. To determine the dynamics of the square lattice system, we perform numerical simulations. We compute the time evolutions for the initial state in the form
$\psi(n,m,0)=(\frac{1}{4\pi^{2}\alpha^{4}})^{1/4}e^{im\pi/2}e^{-(1/4)(m^{2}+n^{2})/\alpha^{2}},$ (50)
in the following four representative cases. (i) To simulate Bloch breathing, the initial state is taken as $\psi(0,m,0)\delta_{n,n_{0}}$ with $\alpha=1$. The system parameters are $\kappa_{0}=J$, $\omega=0$ and $\omega_{0}=0.5J$. (ii) To simulate Bloch oscillation, the initial state is taken as $\psi(n,m,0)$ with $\alpha=1$. The system parameters are $\kappa_{0}=J$, $\omega=0$ and $\omega_{0}=0.5J$. (iii) To simulate superdiffusion, the initial state is taken as $\psi(0,m,0)\delta_{n,n_{0}}$ with $\alpha=1$. The system parameters are $\kappa_{0}=0.25J$ and $\omega=\omega_{0}=0.5J$. (iv) To simulate diffusion, the initial state is taken as $\psi(0,m,0)\delta_{n,n_{0}}$ with $\alpha=1$. The system parameters are $\kappa_{0}=0.25Ji$ and $\omega=\omega_{0}=0.5J$. The evolved state has the form
@@
which can be computed by exact diagonalization. The probability distribution at position $(n,m)$ at time $t$ is defined by the Dirac probabilities
@@
to record the trace of the wavepacket. Here, we use $j\tau$ to denote the discretized time coordinate. It is expected that the distribution of $P(n,m)$  exhibits patterns similar to those in Fig. 3. The numerical results presented in Fig. 4 show that the probability distributions on the square lattice exhibit distinct patterns corresponding to the dynamic behaviors of Bloch breathing, oscillation, diffusion and superdiffusion.
VII Summary

### v3

@@ 5 Dynamical exponents
where $|n\rangle$ denotes the eigenstates of the instantaneous Hamiltonian. The dynamics of $H_{\mathrm{eq}}$ were investigated in Ref. [25], which indicates that the reality of $\kappa_{0}$ plays an important role in the dynamics. The Hamiltonian $H_{\mathrm{eq}}$ can be diagonalized in the form
@@
$\left|\Psi(t)\right\rangle=\sum\limits_{l}a_{l}(t)|l\rangle=\sum\limits_{l}a_{l}(0)\sum_{n}i^{n-l}J_{n-l}(-\kappa_{0}t)|n\rangle,$ (32)
@@
Figure 5: The same plots of $P_{n}(t)/P(t)$ and $\text{In}P(t)$ as those in Fig. 3 are shown for several typical values of system parameters but at a resonance $\mathcal{\omega}_{0}=\omega=1$. Here, (a), (b), and (c) correspond to the cases with $\kappa_{0}=1$, $i$, and $e^{i\pi/4}$, respectively. The middle panels correspond to the probability distributions at moments $t=3T$, $6T$, and $9T$ , respectively. The right panels are the logarithmic plots of total probabilities as functions of time. These results agree with our analytical analysis in the main text. We can see that the patterns in (b) and (c) are almost identical. This inidcates that the imaginary part of hopping plays the important role in the energy-level spreading.
@@
The feature of the Bessel function indicates that there exists a maximum at the edge of $\mathcal{P}_{n}(t)$, which can be regarded as the wave front of the occupied-energy-level spreading. The location of such a wave front can then be determined by the equation
${\frac{\partial\mathcal{P}_{n_{c}}(t)}{\partial t}=0.}$ (36)
Based on the relations of Bessel functions
$J_{n_{c}-1}(x)-J_{n_{c}+1}(x)=2J_{n_{c}}^{\prime}(x),\,\,\,J_{n_{c}+1}\left(x\right)+J_{n_{c}-1}\left(x\right)=2\left|n_{c}/t\right|J_{n_{c}}\left(x\right),$ (37)
we have
${2J_{n_{c}+1}(-t)=2|n_{c}/t|J_{n_{c}}(-t).}$ (38)
For large $n$ and $t$, we have ${J_{n_{c}+1}(-t)\approx J_{n_{c}}(-t)}$, which results in
${|n_{c}|\approx t.}$ (39)
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1$.
(ii) $\kappa_{0}=i$; in parallel, the corresponding amplitude is the Bessel function with imaginary argument
$a_{2n}(t)=(-1)^{n}J_{2n}(-it)=\frac{(-1)^{n}}{2\pi}\int_{-\pi}^{\pi}\cos(2nk)e^{-t\sin k}\mathrm{d}k,$ (40)
@@
$a_{2n}(t)\approx e^{-\frac{(2n)^{2}}{2t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-t\sin k}\mathrm{d}k,$ (42)
@@
$a_{2n+1}(t)\approx e^{-\frac{(2n+1)^{2}}{2t}}\frac{1}{2\pi i}\int_{-\pi}^{\pi}e^{-t\sin k}\mathrm{d}k,$ (43)
@@
$\mathcal{P}_{n}(t)=|a_{n}(t)|^{2}\approx\sqrt{\frac{1}{\pi t}}\mathcal{P}e^{-\frac{n^{2}}{t}},$ (44)
@@
Obviously, the profile of $\mathcal{P}_{n}$ is a Gaussian type, and the total probability grows exponentially as $e^{2t}$ approaches the large $t$ limit. In this situation, the dynamics of occupied-energy-level spreading can be characterized by the full width at half maximum,
@@
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1/2$.
The appendix Appendix presents the dynamical exponent for $\kappa_{0}$ as any complex number. When it is real, $z=1$; when it is not real, $z=1/2$. The non-Hermitian case differs from the Hermitian case primarily due to the imaginary part of $\kappa_{0}$. To verify and demonstrate the above analysis, numerical simulations are performed to investigate the dynamic behavior of dynamic processes at resonance with $\mathcal{\omega}=\mathcal{\omega}_{0}$. For the same initial state as shown in Fig. 3, the plots of $P_{n}(t)$ and $P(t)$ for several typical values of $\kappa_{0}$  are presented in Fig. 5. These numerical results agree with our above analysis in two aspects: (i) At resonance, the dynamics are no longer periodic; (ii) The level spreading exhibits distinct behaviors for Hermitian and non-Hermitian systems.
6 2D simulator via wavepacket dynamics
In this section, we will show how the time evolution of a wavepacket in an engineered time-independent 2D square lattice can simulate our results for Floquet dynamics in a 1D time-dependent lattice. In experiments, single-particle hopping dynamics can be simulated by discretized spatial light transport in an engineered 2D square lattice of evanescently coupled optical waveguides [50]. A 2D lattice can be fabricated by coupled waveguides, by which the temporal evolution of the single-particle probability distribution in the 2D lattice can be visualized by the spatial propagation of the light intensity.
We consider a single-particle Hamiltonian on a square lattice
$H_{\mathrm{2D}}$ $=$ $\kappa_{0}\sum_{n,m=-\infty}^{\infty}\cos\left(qm\right)|n,m\rangle\langle n+1,m|-J\sum_{n,m=-\infty}^{\infty}|n,m\rangle\langle n,m+1|+\mathrm{h.c.}$ (48)
@@
which is an array of coupled 1D Bloch systems with $m$-dependent hopping strength $\kappa_{0}\cos\left(qm\right)$. Fig. 1(b) shows a schematic of the system. According to the above analysis, the system is governed by Bloch dynamics with frequency $\omega_{0}$ for vanishing $J$, regardless of whether $\kappa_{0}$ is real or complex.
Figure 6: For numerical simulations of the dynamics, the trajectories of wavepackets evolving in the in a 2D lattice with four typical configurations were obtained. The traces of the wavepacket $P(n,m)$ for the time evolutions of the four cases are shown in (a1), (b1), (c1), and (d1). The probability distributions $p(n,m,t)$ for the time evolutions of four representative cases at $t=0$, $t=2\pi/J$, $t=4\pi/J$, $t=6\pi/J$ and $t=8\pi/J\,$are shown in (a2), (b2), (c2), and (d2), respectively. The initial states and system parameters are given in the main text. The size of the system is $60\times 30$. The energy scale of the Hamiltonian is taken as $J=1$. The probability distributions on the square lattice exhibit distinct patterns corresponding to the dynamic behaviors of Bloch breathing, oscillation, diffusion and superdiffusion.
@@
$i\frac{\partial}{\partial t}|\psi\rangle=H_{\mathrm{2D}}|\psi\rangle,$ (49)
@@
which resembles that in Ref. [51], where $f\left(k_{x},t\right)$ obeys the following equation:
@@
In the case with $k_{y}\approx\pi/2$, we have $\cos k_{y}\approx k_{y}-\pi/2$. One can construct a solution by the superposition of states $|k_{x},k_{y}\rangle$ with $k_{y}\approx\pi/2$ in the form [52]
$\left|g\left(k_{x},t\right)\right\rangle$ $=$ $(\frac{2\alpha^{2}}{\pi})^{1/4}\int_{-\pi}^{\pi}e^{-\alpha^{2}\left(k_{y}-\pi/2\right)^{2}}e^{it2J\cos k_{y}}f\left(k_{x},t\right)\mathrm{d}k_{y}|k_{x},k_{y}\rangle$ (54)
$\approx$ $(\frac{1}{2\pi\alpha^{2}})^{1/4}\sum_{m}e^{im\pi/2}e^{-(1/4)(m-2Jt)^{2}/\alpha^{2}}f\left(k_{x},t\right)|k_{x},m\rangle,$
which is a traveling Gaussian wavepacket in the $y$ direction. It is presumed that such a solution still holds true by taking $\kappa_{0}\rightarrow\kappa_{0}\cos\left(qm\right)$ when $\kappa_{0}\cos\left(qm\right)$ varies slowly during the extension of the wavepacket in the y direction due to the locality of the solution. In this sense, the traveling wavepacket experiences an adiabatic process, after which the independence of the dynamics in both directions is maintained. Obviously, a small value of $q$ should be needed. On the other hand, the value of $\omega_{0}$ must match $q$ when the system is employed to simulate the dynamics at the resonance. To this end, a small value of $|\kappa_{0}|$ should be taken to maintain the Stark ladder.
Actually, under the conditions $|\kappa_{0}|\ll J$ and $k_{y}\approx\pi/2$, such a 2D system is equivalent to a coupled waveguide array due to the linear dispersion in the $y$ direction, in which the tunneling rate between neighboring waveguides is $m$ dependent. The speed of the propagating wavepacket along $m$ is $2J$. To simulate the dynamics along $x$, one can simply take $q=\omega/\left(2J\right)$, regarding $m$ as the temporal coordinate.
To simulate the time-dependent evolution of a 1D system (1) in a 2D lattice system (48), we design specific initial states
$\psi(n,m,0)=x(n)y(m),$ (55)
where $x(n)$ is taken as the initial state of the 1D system and $y(m)=({4\pi^{2}})^{-1/4}e^{im\pi/2}e^{-m^{2}/4}$ represents a Gaussian wave packet with momentum ${\pi}/{2}$ in the y-direction, serving as an alternative to the time axis. At first, we verify this scheme by considering the Bloch oscillation in Hermitian system. (i) To simulate Bloch breathing, $x(n)$ is taken as $\delta_{n,n_{0}}$. The system parameters are $\kappa_{0}=J$, $\omega=0$ and $\omega_{0}=0.5J$. (ii) To simulate Bloch oscillation of a wavepacket, $x(n)$ is taken as $e^{-{(n-n_{0})^{2}}/4}$. The system parameters are $\kappa_{0}=J$, $\omega=0$, and $\omega_{0}=0.5J$. Second, we consider the simulations for the dynamics, which are presented in Fig. 4(b1) and (b2). (iii) To simulate superdiffusion, $x(n)$ is taken as $\delta_{n,n_{0}}$. The system parameters are $\kappa_{0}=0.25J$ and $\omega=\omega_{0}=0.5J$. (iv) To simulate diffusion, $x(n)$ is taken as $\delta_{n,n_{0}}$. The system parameters are $\kappa_{0}=0.25Ji$ and $\omega=\omega_{0}=0.5J$. For greater clarity, the specific parameters are shown in Tab. 1. The evolved state has the form
@@
which can be computed by exact diagonalization.
Table 1: List of the parameters of the four driven Hamiltonians and initial states, where $y(m)=({4\pi^{2}})^{-1/4}e^{im\pi/2}e^{-m^{2}/4}$ represents a Gaussian wave packet with momentum ${\pi}/{2}$ in the y-direction, moving at a constant velocity $v=2J$ in the y-direction. Therefore, the distance the wave packet travels in the y-direction represents time.
       $\kappa_{0}$        $\omega$        $\omega_{0}$        Initial state
[…]
The probability distribution at position $(n,m)$ at time $t$ is defined by the probabilities
@@
to record the trace of the wavepacket. Here, we use $j\tau$ to denote the discretized time coordinate. It is expected that the distribution of $P(n,m)$  exhibits patterns similar to those in Fig. 5. The numerical results presented in Fig. 6 show that the probability distributions on the square lattice exhibit distinct patterns corresponding to the dynamic behaviors of Bloch breathing, oscillation, superdiffusion and diffusion.
7 Conclusion
@@ Acknowledgements
Appendix
In this appendix Appendix, we derive the dynamical exponent for the case in which $\kappa_{0}$ is an arbitrary complex number. When $\kappa_{0}$ is a real number, i.e., $\text{Im}(\kappa_{0})=0$, using the same Bessel function estimation method as in the main text, one can obtain the dynamical exponent of occupied-energy-level spreading $z=1$, and the total probability $\mathcal{P}$ remains constant at 1.
When $\text{Im}(\kappa_{0})\neq 0$, we have
$a_{2n}(t)=(-1)^{n}J_{2n}(-\kappa_{0}t)=\frac{(-1)^{n}}{2\pi}\int_{-\pi}^{\pi}\cos(2nk)e^{i\kappa_{0}t\sin k}\mathrm{d}k,$ (59)
and
$a_{2n+1}(t)=i(-1)^{n}J_{2n+1}(-\kappa_{0}t)=\frac{(-1)^{n+1}}{2\pi i}\int_{-\pi}^{\pi}\sin[(2n+1)k]e^{i\kappa_{0}t\sin k}\mathrm{d}k.$ (60)
On a long time scale, the main contribution to the integral comes from the integrand around $k=-\pi/2$(or $\pi/2$, depends on the sign of $\text{Im}(\kappa_{0})$). Then, we have
$a_{2n}(t)$$\approx$ ${e^{-\frac{(2n)^{2}}{2i\kappa_{0}t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})<0}$ (61)
$\approx$ ${e^{\frac{(2n)^{2}}{2i\kappa_{0}t}}\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})>0}$ (62)
and
$a_{2n+1}(t)$$\approx$ ${e^{-\frac{(2n+1)^{2}}{2i\kappa_{0}t}}\frac{i}{2\pi}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})<0}$ (63)
$\approx$ ${e^{\frac{(2n+1)^{2}}{2i\kappa_{0}t}}\frac{1}{2\pi i}\int_{-\pi}^{\pi}e^{i\kappa_{0}t\sin k}\mathrm{d}k,\,\text{Im}(\kappa_{0})>0}$ (64)
$\mathcal{P}_{n}(t)=|a_{n}(t)|^{2}\approx\sqrt{\frac{|\text{Im}(\kappa_{0})|}{\pi|\kappa_{0}|^{2}t}}\mathcal{P}e^{-\frac{|\text{Im}(\kappa_{0})|n^{2}}{|\kappa_{0}|^{2}t}},$ (65)
with the total probability
$\mathcal{P}=\sum_{n}\mathcal{P}_{n}=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-2\text{Im}(\kappa_{0})t\sin k}\mathrm{d}k.$ (66)
Obviously, the profile of $\mathcal{P}_{n}$ is a Gaussian type, and the total probability grows exponentially as $e^{2|\text{Im}(\kappa_{0})|t}$ approaches the large $t$ limit. In this situation, the dynamics of occupied-energy-level spreading can be characterized by the full width at half maximum,
$\mathcal{P}_{n_{c}}(t)=\frac{1}{2}\mathcal{P}_{0}(t),$ (67)
which results in
$|n_{c}|\approx\sqrt{t}.$ (68)
Then, we conclude that the dynamical exponent of occupied-energy-level spreading is $z=1/2$.